an eighteen gauge copper wire has a nominal diameter of 1.02mn.this wire carries a constant current of 1.67a to a 200w lamp.the density of free electrons is 8.5×10^28 electrons per cubic metre find magnitude of the current density and the drift velocity


1
Expert's answer
2021-05-23T16:33:13-0400
"j=\frac{1.67}{0.25\pi(1.02\cdot10^{-3})^2}=2.04\cdot10^{6}\frac{A}{m^2}\\v=\frac{2.04\cdot10^{6}}{(8.5\cdot10^{28})(1.6\cdot10^{-19})}=1.5\cdot10^{-4}\frac{m}{s}"